Suppose utility function is given by U(x,y) = x0.3y0.7 subject to budget constraint I= Pxx + Pyy. Further suppose that the prices for those goods given by Px=1 Py=2 & I=2. Derive demand function of good x and good y.


1
Expert's answer
2020-07-08T17:47:59-0400
"\frac {\delta U}{\delta x}=0.3(\frac{y}{x})^{0.7}"


"\frac {\delta U}{\delta y}=0.7(\frac{x}{y})^{0.3}"


"\frac {\frac {\delta U}{\delta x}}{\frac {\delta U}{\delta y}}=\frac {p_x}{p_y}"


"\frac {x}{y}=\frac{6}{7}"

"x=\frac {6}{7}y"

"\frac {6}{7}y+2y=I"


Thus, the demand function for a set of two products has the form


"y=\frac{7}{20}I"


"x=\frac{3}{10}I"